OpenStudy (anonymous):

-20= -7 - 4 (x=2) + 5(3x - 1)

OpenStudy (anonymous):

K i got this

OpenStudy (anonymous):

how do I make to where I just get x= whatever it is

OpenStudy (anonymous):

I am on it give me a sec

OpenStudy (anonymous):

ok thank you

OpenStudy (anonymous):

X=* what is that

OpenStudy (anonymous):

That has to be a mistake right?

OpenStudy (anonymous):

I meant (x+2) not (x=2)

OpenStudy (anonymous):

Alright then it can be done

OpenStudy (anonymous):

ok

OpenStudy (jackellyn):

@samissmurfette , you need to distribute and combine like terms, then you can solve for x

OpenStudy (anonymous):

I don't understand how to distribute and combine the terms. ive tried and I cant seem to do it.

OpenStudy (anonymous):

Well this is what i am getting x=0

OpenStudy (anonymous):

x=0? ok but how do you get to that?

OpenStudy (jackellyn):

-20= -7 - 4 (x+2) + 5(3x - 1) To distribute, you multiply the number in front of the parenthesis to the numbers/variable inside. Example: 2(x-4) becomes 2(x)-2(4)=2x-8

OpenStudy (anonymous):

-20=-7-4(x+2)+5(3x-1) -20=-7-4x-8+15x-5 X=0

OpenStudy (jackellyn):

So -20=-7-4x-8+15x-5 to combine like terms, you get all the numbers together that dont have variables and then all the numbers that do have variables together. -20+7+8+5=-4x+15x

OpenStudy (jackellyn):

The left hand side would add up to 0. 0=11x 0=x

OpenStudy (anonymous):

^i confirm

OpenStudy (jackellyn):

Does that make sense?

OpenStudy (anonymous):

sorta. im kinda slow. im sure if I looked at it long enough I would understand. thank you for your help

OpenStudy (jackellyn):

@samissmurfette , you just need practice! :)

OpenStudy (anonymous):

yeah I do. haha.

OpenStudy (jackellyn):

Anything I can help clear up?

OpenStudy (anonymous):

no. I got it now. its just we did it differently in class.

OpenStudy (jackellyn):

Show me how you would do it in your class

OpenStudy (anonymous):

well we wouldn't add all the terms together. we didn't add -20 in there. it would only be 8 7 and 5 and then we would subtract 8 7 and 5 from -20 and then we divide

OpenStudy (jackellyn):

Right, it would give the same answer in the end

OpenStudy (anonymous):

yeah you just made it shorter for me. not as much work @jackellyn